OpenStudy (anonymous):

HELP PLEASE Stan evaluated a function with a value and got 3 as the result. He wrote the entire process as a sentence: The function , when evaluated with , gives the value 3.

OpenStudy (anonymous):

okay gimmie a second

OpenStudy (anonymous):

so u have to make it an actual problem

OpenStudy (anonymous):

yes

OpenStudy (mertsj):

The function, when evaluated with an unknown number, gives a value of 3.

OpenStudy (anonymous):

does it matter how i do the fucntion

myininaya (myininaya):

I don't get it.

myininaya (myininaya):

Is there something missing?

myininaya (myininaya):

Is there a question?

OpenStudy (anonymous):

no theres not u have to make the sentence an actual function

myininaya (myininaya):

What's the question?

myininaya (myininaya):

I don't see one.

OpenStudy (anonymous):

function answers are f(x)=x^2+2x and f(x)=x^2-3x

myininaya (myininaya):

oh you are suppose to make up a function? f(x)=3 or f(x)=x could also work

OpenStudy (mertsj):

f(x)=x+3 When evaluated with x = 0, the result is 3 since f(0)-0+3=3

myininaya (myininaya):

I like f(x)=3 because it is pretty cute because any number you plug into f will give you 3